OpenStudy (anonymous):

Solve the system using substitution 3x+2y=-14 and y=x+3

OpenStudy (anonymous):

3x-2y=14 x-y=-3 3x-2y=-14 (-2)(x-y=-3) 3x-2y=-14 -2x+2y=6 x=-8 Then plug in x to one of the equations

OpenStudy (mathstudent55):

3x + 2y = -14 y = x + 3 Substitution means you solve one equation for one unknown and substitute it into the other equation. In this case, the second equation is already solved for y. To substitute, substitute y in the first equation by x + 3. 3x + 2(x + 3) = -14 Now we solve this equation for x: 3x + 2x + 6 = -14 5x = -20 x = -4 Now that we know x, we substitute x into one of the two original equations to find y. Let's use the second equation: y = -4 + 3 y = -1 The solution is: x = -4, y = -1

OpenStudy (anonymous):

oops lol i didny do substitution.. good job @mathstudent55

OpenStudy (mathstudent55):

@Daniellee The answers should be the same no matter which method you use. You tried to solve the system by addition, but you inadvertently changed the first equation to a different equation. That's why your answer is incorrect. y = x + 3 is the same as x - y = -3 3x + 2y = -14 is not the same as 3x - 2y = 14